# Conditional Generative Representation for Black-Box Optimization with Implicit Constraints

### v3

@@ 5 Conclusion
In this paper, we introduce and study a new category of optimization problems termed Implicit-Constrained Black-Box Optimization (ICBBO) where constraints are implicit but can be easily verified. This problem is well motivated from real-world police redistricting problems where districting plans are constrained to socio-economic or political considerations but feasible plans can be easily verified by policymakers. This new ICBBO problem poses more challenges than conventional black-box optimization where existing work cannot be readily applied.
@@
To address this problem, we develop the CageBO algorithm and its core idea is to learn a conditional generative representation of feasible decisions to effectively overcome the complications posed by implicit constraints. In this way, our method can not only learn a good representation of original space but also generate feasible samples that can be used later. Moreover, our algorithm facilitates Bayesian optimization to search in an unconstrained and more compact latent space. Our theoretical analysis shows that CageBO can find global optimal solution by expectation and empirical studies show its effectiveness in real-world police redistricting case in Atlanta, Georgia. More broadly, our method can also accommodate traditional constraints by first generating samples that satisfy these constraints and then employing our algorithm for optimization, which yields another potential use case for our framework.
Appendix A Details of theoretical analysis
In this section, we restate our main theorem and show its complete proof afterwards.
Theorem 3 (Restatement of Theorem 1).
@@ Appendix A Details of theoretical analysis
Our proof starts from the bounding the error term incurred in post-decoder. Let $S$ denote the set of $n$ data points sampled i.i.d. from domain $\mathcal{X}$ and $\epsilon$ denote the expected distance between any data point $x$ and its nearest neighbor $x^{\prime}$ in $\mathcal{X}$, i.e.,
@@
Recall that $\mathcal{X}\subseteq[0,1]^{d}$ and we discretize it in each dimension using distance $\varepsilon$ and we get $r=(1/\varepsilon)^{d}$ small boxes. Each small box $C_{i}\forall i\in 1,...,r$ is a covering set of the domain. For any two data points in the same box we have $\|x_{1}-x_{2}\|\leq\sqrt{d}\varepsilon$, otherwise, $\|x_{1}-x_{2}\|\leq\sqrt{d}$. Therefore,
@@
Here are two events: with probability $C_{f}$, $\hat{z}_{t}$ suggested by GP-LCB is feasible and post-decoder is not needed and with probability $1-C_{f}$, $\hat{z}_{t}$ needs to be post-decoded. Thus,
$\mathbb{E}[R_{T}]$ $=C_{f}\mathbb{E}\left[\sum_{i=1}^{T}g(\hat{z}_{t})-g(z_{*})\right]+(1-C_{f})\mathbb{E}\left[\sum_{i=1}^{T}g(\hat{z}_{t})-g(z_{*})+g(z_{t})-g(\hat{z}_{t})\right],$ (15)
@@
Technical lemmas. Our proof relies on the following two lemmas and full proof is shown in Appendix A.
Lemma 4 (Regret bound of GP-LCB (Theorem 1 of [39])).
@@
Lemma 5 (Expected minimum distance (Lemma 19.2 of [35])).
@@
Appendix B Implementation details of the conditional generative model
@@
$\mathcal{L}_{\mathrm{ELBO}}=w(c)\underset{q(z|x,c)}{\mathbb{E}}\left[\log p(x|z,c)\right]-\eta\mathrm{D}_{\mathrm{KL}}\left(q(z|x,c)\|p(z|c)\right).$ (28)
In practice, we approximate the conditional distribution $q(z|x,c)$ and $p(x|z,c)$ by building the ${\mathrm{encoder}}_{\phi}$ and ${\mathrm{decoder}}_{\theta}$ neural networks, and denote the approximated conditional distribution to be $q_{\phi}(z|x,c)$ and $p_{\theta}(x|z,c)$ respectively. In our formulation, the prior of the latent variable is influenced by the feasibility $c$; however, this constraint can be easily relaxed to make the latent variables statistically independent of their labels, adopting a standard Gaussian distribution $\mathcal{N}(0,I)$. Both $q(z|x,c)$ and $p(x|z,c)$ are typically modeled as Gaussian distributions to facilitate closed-form KL divergence computation. Specifically, we introduce generator layers $l(\epsilon,x,c)$ and $l(\epsilon)$ to represent $q(z|x,c)$ and $p(z)$, respectively, transforming the random variable $\epsilon\sim\mathcal{N}(0,I)$ using the reparametrization trick [37]. The log-likelihood of the first term is realized as the reconstruction loss, computed with the training data.
Appendix C Ablation study
@@
Appendix D Limitation and societal impact
It is worth mentioning that this work shares similar limitations to prior research on latent-space Bayesian optimization with VAEs, and it could have comparable societal impacts to the “predictive policing" systems deployed several years ago, which sparked significant controversy. We acknowledge the importance of addressing these concerns, particularly in the context of police redistricting based on historical data, which might inadvertently prioritize areas with historically higher incident rates.
To mitigate such biases, our approach includes debiasing the historical data using machine learning techniques to ensure fairness in districting plans. Additionally, our CageBO algorithm is designed to accommodate implicit constraints aimed at promoting equitable districting, as suggested by practitioners. These measures align with our commitment to consider the broader impact of our work and the ethical implications of implementing machine learning solutions in public policymaking.

### v4

@@ Conclusion
In this paper, we introduce and explore a new category of optimization problems in public policy, termed Implicit-Constrained Black-Box Optimization (ICBBO), where constraints are implicit but can be easily verified. This problem is well motivated by real-world police redistricting problems where districting plans are constrained to socio-economic or political considerations but feasible plans can be easily verified by policymakers. This new ICBBO problem poses more challenges than conventional black-box optimization where existing work cannot be readily applied.
@@
To address this problem, we develop the CageBO algorithm and its core idea is to learn a conditional generative representation of feasible decisions to effectively overcome the complications posed by implicit constraints. In this way, our method can not only learn a good representation of the original space but also generate feasible samples that can be used later. Our theoretical analysis confirms that CageBO can, on expectation, identify the global optimal solution. This approach was empirically validated through a case study on police redistricting in Atlanta, Georgia, where our algorithm demonstrated superior performance in optimizing districting plans. This case study highlights the potential social impact of CageBO, offering a tool that can contribute to more equitable and efficient public policy decisions. Beyond this application, our method is adaptable enough to incorporate traditional constraints by generating compliant samples before optimization, extending its utility across a wide range of policy-making scenarios.
Appendix A Appendix
Details of Theoretical Analysis
In this section, we restate our main theorem and show its complete proof afterward. In this section, we restate our main theorem and provide its complete proof. We begin by formally stating the two key assumptions that will be utilized. The first assumption pertains to the properties of the encoder.
Assumption 3.
The distance between any two points in $\mathcal{Z}$ can be upper bounded by their distance in $\mathcal{X}$, i.e.,
$\|\mathtt{enc}(x)-\mathtt{enc}(x^{\prime})\|_{2}\leq C_{p}\|x-x^{\prime}\|_{2},\quad\forall x,x^{\prime}\in\mathcal{X}.$ (8)
where $C_{p}$ is a universal constant.
The next assumption is the standard Gaussian process assumption for Bayesian optimization (Srinivas et al. 2010).
Assumption 4.
Function $g:\mathcal{Z}\mapsto\mathbbm{R}$ is drawn from some Gaussian Process and it is $C_{g}$-Lipschitz continuous, i.e.,
$|g(z)-g(z^{\prime})|\leq C_{g}\|z-z^{\prime}\|_{2},\quad\forall z,z^{\prime}\in\mathcal{Z},$ (9)
where $C_{g}$ is a universal constant.
Theorem (Restatement of Theorem 1).
@@ Appendix A Appendix
where $\gamma_{T}$ is the maximum information gain, depending on the choice of kernel used in the algorithm and $n$ is the number of initial observation data points.
@@
Our proof starts from the bounding the error term incurred in the post-decoder. Let $S$ denote the set of $n$ data points sampled i.i.d. from domain $\mathcal{X}$ and $\epsilon$ denote the expected distance between any data point $x$ and its nearest neighbor $x^{\prime}$ in $\mathcal{X}$, i.e.,
@@
Recall that $\mathcal{X}\subseteq[0,1]^{d}$ and we discretize it in each dimension using distance $\varepsilon$ and we get $r=(1/\varepsilon)^{d}$ small boxes. Each small box $C_{i}\forall i\in 1,...,r$ is a covering set of the domain. For any two data points in the same box, we have $\|x_{1}-x_{2}\|\leq\sqrt{d}\varepsilon$, otherwise, $\|x_{1}-x_{2}\|\leq\sqrt{d}$. Therefore,
@@
Here are two events: with probability $C_{f}$, $\hat{z}_{t}$ suggested by GP-LCB is feasible and post-decoder is not needed, and with probability $1-C_{f}$, $\hat{z}_{t}$ needs to be post-decoded. Thus,
$\mathbb{E}[R_{T}]$ $=C_{f}\mathbb{E}\left[\sum_{i=1}^{T}g(\hat{z}_{t})-g(z_{*})\right]$ (17)
$+(1-C_{f})\mathbb{E}\left[\sum_{i=1}^{T}g(\hat{z}_{t})-g(z_{*})+g(z_{t})-g(\hat{z}_{t})\right],$ (18)
@@
Technical lemmas. Our proof relies on the following two lemmas:
Lemma 5 (Regret bound of GP-LCB (Theorem 1 of Srinivas et al. (2010))).
@@
Lemma 6 (Expected minimum distance (Lemma 19.2 of Shalev-Shwartz and Ben-David (2014))).
@@
Implementation Details
@@
$\mathcal{L}_{\mathrm{ELBO}}$ $=w(c)\underset{q(z|x,c)}{\mathbb{E}}\left[\log p(x|z,c)\right]$
$\quad-\eta\mathrm{D}_{\mathrm{KL}}\left(q(z|x,c)\|p(z|c)\right).$ (31)
In practice, we approximate the conditional distribution $q(z|x,c)$ and $p(x|z,c)$ by building the ${\mathrm{encoder}}_{\phi}$ and ${\mathrm{decoder}}_{\theta}$ neural networks, and denote the approximated conditional distribution to be $q_{\phi}(z|x,c)$ and $p_{\theta}(x|z,c)$ respectively. In our formulation, the prior of the latent variable is influenced by the feasibility $c$; however, this constraint can be easily relaxed to make the latent variables statistically independent of their labels, adopting a standard Gaussian distribution $\mathcal{N}(0,I)$. Both $q(z|x,c)$ and $p(x|z,c)$ are typically modeled as Gaussian distributions to facilitate closed-form KL divergence computation. Specifically, we introduce generator layers $l(\epsilon,x,c)$ and $l(\epsilon)$ to represent $q(z|x,c)$ and $p(z)$, respectively, transforming the random variable $\epsilon\sim\mathcal{N}(0,I)$ using the reparametrization trick (Sohl-Dickstein et al. 2015). The log-likelihood of the first term is realized as the reconstruction loss, computed with the training data.
Ablation Study
@@
Debiasing for Fairness
It is worth mentioning that this work shares the same concerns with prior research on latent-space Bayesian optimization using VAEs (Notin, Hernández-Lobato, and Gal 2021), where the generated latent variables often fail to consistently represent the underlying data distributions. Such limitations could lead to social impacts similar to those observed with predictive policing systems (Meijer and Wessels 2019), which have faced significant controversy due to their reinforcement of existing biases. Specifically, redistricting to minimize workload variance based on historical 911 calls can inadvertently justify the concentration of policing efforts in districts with a history of higher incident rates. We recognize the critical importance of addressing these ethical concerns, particularly in the context of police redistricting based on historical data, which might inadvertently prioritize areas with historically higher incident rates.
To mitigate these biases, our approach involves debiasing historical data using established machine learning techniques to ensure fairness in districting plans (Zhu and Xie 2022; Jiang and Xie 2024). Furthermore, our CageBO algorithm is designed to accommodate implicit constraints aimed at promoting equitable districting, as suggested by practitioners. These measures align with our commitment to consider the broader impact of our work and the ethical implications of implementing machine learning solutions in public policymaking.